OpenStudy (anonymous):

PLEASE HELP MEEE Miranda bought some apples and oranges at the store for fruit salad. Owen bought three times as many apples and twice as many oranges. Miranda has 5 fruits in her bag, but Owen has 12 fruits in his bag. How many apples did Miranda purchase? 1 2 3 5

OpenStudy (jack1):

Miranda: x +y = 5 owen: 3x +2y = 12 x = apples y = oranges ...now solve simultaneous equation... i'd recommend substitution

OpenStudy (anonymous):

How do I do substitution?

OpenStudy (anonymous):

I did this in like 8th grade but I cant remember how, can u restart my brain by giving me a push in the right direction?

OpenStudy (jack1):

Miranda: x +y = 5 Equation 1 Owen: 3x +2y = 12 Equation 2 rearrange equation 1 to get y = 5 - x sub the new equation into equation 2 3x + 2y = 12 3x + 2(5 -x) = 12 so x = ...?

OpenStudy (anonymous):

Ok.. so far to figure it out I subtracted 3 and subtracted it from the 12... and got x+2(5-x)=9

OpenStudy (anonymous):

Is that correct so far?

OpenStudy (jack1):

no

OpenStudy (anonymous):

..Ok. Let me try something else.

OpenStudy (jack1):

3x + 2y = 12 3x + 2(5-x) = 12 3x + 10 -2x = 12 3x - 2x = 12 - 10 x = 12-10 x = 2 so mirinda originally has 2 apples and as she starts out with 5 fruit, she also has 3 oranges

OpenStudy (anonymous):

Ok so u distribute the 2 with whats in the parenthesis first?

OpenStudy (jack1):

maybe i described it poorly at the start x = number of apples Mirinda has y = number of oranges Mirinda has and yeah, use BODMAS to do the correct order of operations

OpenStudy (anonymous):

BODMAS? I only know of PEMDAS?

OpenStudy (jack1):

ok use that then, im in Aus, so they teach it differently down here, sorry ;)

OpenStudy (anonymous):

Oh thank you so much! I just didn't do my homework for my next class and im trying to hustle through it but I don't want to get them wrong. Thank u.. I took notes to try for next time.

OpenStudy (jack1):

so yeah: expand out the brackets add/subtract like terms divide if necessary to isolate the term u want

OpenStudy (jack1):

all good, always welcome

OpenStudy (anonymous):

let the number of apple and oranges bought by miranda be x and y respectively,Therefore, x+y=5(Miranda) and, 3x+2y=12(owen) y=5-x....Substituting this in Owen's equation.3x+2(5-x)=12 will be,3x+10-2x=12 Group like terms together x=12-10 x=2 and y=5-2=3 Therefore,Miranda has 2 Apples